Fermi National Laboratory


Accelerator Update for January 7, 2005 - January 10, 2005

Friday January 7
The day shift began with the Recycler stashing antiprotons, and with MiniBooNE and the Meson experiments, MTest and MCenter, taking beam.

Operations put the Tevatron (TeV) into shot setup at 8:23 AM.

Operations established store 3908 at 11:28 AM with an initial luminosity 76.81E30. The Antiproton Source (Pbar) resumed stacking.

At 2:08 PM, the TeV suffered an abort, its RF tripped off and the store was lost. Experts investigated and found that a phase detector had lost power, which inhibited the anode program signal.

Operators made the Collider Detector at Fermilab (CDF) ready for access. Among other jobs, cryo system technicians will work on the TeV sector A4 wet engine.

CDF came out of access at 5:25 PM.

Operations put the TeV through a dry squeeze at 5:42 PM.

TeV experts began a study period at 6:58 PM.

Operations established beam to MCenter at 8:38 PM. MTest shut down at 10:22 PM.

Saturday January 8
The midnight shift began with Pbar stacking, with Operations monitoring the stash, with the TeV conducting studies, and with MiniBooNE and MCenter taking beam.

At 12:17 AM, the TeV suffered a quench on the C4 and D0 low beta quadrupole magnets. The problem was later discovered to be bad table parameters.

At 2:27 AM, the MiniBooNE horn tripped off. The horn continued to trip until operators discovered the trips were associated with the new timeline started for slipstacking and beam to Switchyard (SY). When operators reverted to the former timeline, the trips stopped.

Operations put the TeV into shot setup at 3:32 AM.

At 3:40 AM, the TeV suffered a sector F3 quench. Operators think the quench was due to a timeline problem. Pbar resumed stacking.

At 5:30 AM, the TeV quenched due to a problem with a power supply (T:QDD1). Later it was found that the power supply's quench card was dirty.

Operations put the TeV into shot setup at 11:41 AM.

Operations established store 3915 at 2:39 PM with an initial luminosity of 67.25E30. Pbar resumed stacking

Operations began a Recycler shot setup at 7:35 PM. After the transfers, Pbar resumed stacking at 9:07 PM.

Sunday January 9
The midnight shift began with Operations monitoring store 3915 and the stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 5:45 AM, two Booster RF stations (BRF7 & 8) tripped off due to a vacuum burst. The stations reset okay.

Operations terminated store 3915 at 5 PM. Earlier, the TeV sector F17 kicker had tripped off. Operators tried to reset it after terminating the store, but it didn't work. Operators prepared the TeV for an access into F sector. An F sector access shuts off beam to all experiments. CDF experimenters also made an access into the TeV during this time.

Everyone was out of the TeV by 7:28 PM.

Operations had beam to MiniBooNE by 8:50 PM.

Pbar resumed stacking at 9:24 PM.

Operations put the TeV into shot setup at 11:13 PM.

Monday January 10
The midnight shift began with the TeV in shot setup and with MiniBooNE taking beam. Beam for stacking and to SY is always disabled during shot setup.

Operations established store 3917 at 2:41 AM with an initial luminosity of 75.68E30. Pbar resumed stacking.

Operations resumed sending beam to MCenter at 3:39 AM. MCenter beam was delayed due to a problem with a power supply (S:HPP3DS).

MCenter shut off for the day at 7:54 AM.

The Plan for Monday
The plan for this morning is to stash antiprotons in the Recycler. The plan for later this afternoon, and from now on whenever possible, is to establish a mixed Pbar mode store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/10/2005   email Fermilab